The life of a compressor manufactured by a company is known to be 200 months on an
average following an exponential distribution .Find the probability that the life of a
compressor is i) less than 200 months ii) between 100 months to 25 years.
1
Expert's answer
2020-04-21T14:41:40-0400

let X=The life of a compressor manufactured by a company 

"\mu=200months\\
\lambda=\frac{1}{\mu}=\frac{1}{200}=0.005"

here "X\thicksim exp(0.005)"

"f(x)= \begin{cases}
 0.005e^{(-0.005x)} &\text{if } x\ge0 \\
 0 &\text{if } x<0
\end{cases}"


and


"F(x)= \begin{cases}
 1-e^{(-0.005x)} &\text{if } x\ge0 \\
 0 &\text{if } x<0
\end{cases}"


i)

"P(X<200)=F(200)\\
P(X<200)=1-e^{(-0.005*200)}\\
P(X<200)=0.6321"

probability that the life of a compressor is less than 200 months=0.6321

ii)

"P(100<X<300)=F(300)-F(100)\\
P(100<X<300)=(1-e^{(-0.005*300)})\\
\hspace{12 em}-(1-e^{(-0.005*100)})\\
P(100<X<300)=0.3834"

probability that the life of a compressor is between 100 months to 25 years=0.3834